### v1

# VF-Plan: Bridging the Art Gallery Problem and Static LiDAR Scanning with Visibility Field Optimization

Abstract
Viewpoint planning is crucial for 3D data collection and autonomous navigation, yet existing methods often miss key optimization objectives for static LiDAR, resulting in suboptimal network designs. The Viewpoint Planning Problem (VPP), which builds upon the Art Gallery Problem (AGP), requires not only full coverage but also robust registrability and connectivity under limited sensor views. We introduce a greedy optimization algorithm that tackles these VPP and AGP challenges through a novel Visibility Field (VF) approach. The VF captures visibility characteristics unique to static LiDAR, enabling a reduction from 2D to 1D by focusing on medial axis and joints. This leads to a minimal, fully connected viewpoint network with comprehensive coverage and minimal redundancy. Experiments across diverse environments show that our method achieves high efficiency and scalability, matching or surpassing expert designs. Compared to state-of-the-art methods, our approach achieves comparable viewpoint counts (VC) while reducing Weighted Average Path Length (WAPL) by approximately 95%, indicating a much more compact and connected network. Dataset and source code will be released upon acceptance.
keywords:
Viewpoint Optimization , Terrestrial Laser Scaning , Visibility Field , Medial axis transform , Greedy searching , BIM
[…]
[label1]organization=School of Computer Science and Artificial Intelligence, Wuhan University of Technology,addressline=122 Luoshi Rd, city=Wuhan, postcode=430070, state=Hubei, country=China
[…]
[label2]organization=College of Computer Science and Technology, Zhejiang University, addressline=866 Yuhangtang Rd, city=Hangzhou, postcode=310058, state=Zhejiang, country=China
[…]
[label3]organization=The Advanced Laser Technology Laboratory of Anhui Province, addressline= 850 Huangshan Rd, city=Hefei, postcode=230031, state=Anhui, country=China
{graphicalabstract}
Viewpoint optimization for static LiDAR Scanning. (a) The input floor plan; (b) Optimized viewpoints indicated by green dots, with the Visibility Field displayed as the background image; (c) and (d) Simulated LiDAR points from optimal viewpoints, shown in 2D and 3D respectively, with each scan distinguished by a unique color.
{highlights}
Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the Art Gallery Problem, achieving efficient coverage with minimal viewpoints and robust network connectivity.
Visibility Field for Dimensionality Reduction: We introduce a continuous Visibility Field tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
Greedy Algorithm for Efficient Coverage: Our greedy Algorithm utilizes the VF’s structural insights to construct a minimal, fully connected viewpoint network, ensuring both robust coverage and network connectivity.
1 Introduction
The Viewpoint Planning Problem (VPP) has emerged as a critical research spot due to the growing use of LiDAR and camera sensors in applications such as 3D reconstruction, autonomous navigation, and Scan-to-BIM applications [41, 30]. VPP focuses on determining the optimal placements and configurations of sensors to achieve specific reconstruction goals, such as maximizing observation completeness and minimizing occlusions, while balancing efficiency and data quality. Both model-free and model-based algorithms for viewpoint and path planning have gained significant attention due to their applicability across various industries, including robotics, construction, and surveillance [31].
Conceptually, VPP is rooted in the Art Gallery Problem (AGP), a classic challenge in computational geometry that seeks the minimal number of “guards” required to cover a given area, typically a polygon, with full 360° visibility [25]. AGP is often simplified by constraining guard placements to polygon vertices; however, this simplification does not capture the true computational complexity of the problem. In its generalized form, AGP is $\exists\mathbb{R}$-complete, meaning it is at least as hard as any NP-complete problem, with no known efficient solutions [11, 1]. This NP-hard nature has inspired substantial research interest, particularly in constrained versions that make idealized assumptions, such as complete visibility and static scenes [11, 41].
Figure 1: Viewpoint optimization for static LiDAR Scanning. (a) The input floor plan; (b) Optimized viewpoints indicated by green dots, with the Visibility Field displayed as the background image; (c) and (d) Simulated LiDAR points from optimal viewpoints, shown in 2D and 3D respectively, with each scan distinguished by a unique color.
VPP, however, introduces additional complexities beyond those of AGP. Unlike static guards, sensors in VPP must maintain overlapping fields of view for accurate data registration, accommodate limited and directional visibility ranges, and meet high accuracy requirements for 3D scene reconstruction. These real-world demands, especially in cluttered or partially obstructed environments, render VPP a more challenging NP-hard problem, requiring innovative optimization strategies beyond those used for AGP [41, 1].
Our work addresses these challenges by introducing a novel greedy approach based on the concept of a Visibility Field (VF), specifically designed for static LiDAR applications. The VF captures the unique visibility characteristics of LiDAR, including constraints on range and incident angles, enabling a significant reduction in computational complexity. By focusing on critical points such as medial axis and joints within the VF, we effectively reduce the optimization space from 2D to 1D, facilitating the construction of a minimal, fully connected viewpoint network that achieves comprehensive coverage with reduced redundancy. Our main contributions are as follows:
  1. 1.
    Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the Art Gallery Problem, achieving efficient coverage with minimal viewpoints and robust network connectivity.
  2. 2.
    Visibility Field for Dimensionality Reduction: We introduce a continuous Visibility Field tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
  3. 3.
    Greedy Algorithm for Efficient Coverage: Our greedy algorithm utilizes the VF’s structural insights to construct a minimal, fully connected viewpoint network, ensuring both robust coverage and network connectivity.
2 Related Work
This section reviews key advancements in sensor network planning, covering both static and mobile sensing, with a focus on recent works in LiDAR and computer vision area. These serve as the foundation for the viewpoint planning method proposed here, addressing gaps in efficient coverage, connectivity, and optimization complexity.
2.1 Station-Based Scanning
Sensor network design shares similarities with the Planning for Scanning (P4S) problem, observed in applications like surveillance cameras [32, 55, 16], directional sensors [15], 5G base stations [42, 45], stealth game [49], and station-based LiDAR scanning [2]. For LiDAR, the goal is to cover the entire scene with comprehensive wall scanning while ensuring data quality criteria such as completeness and accuracy [29, 2].
Advancements in LiDAR technology have led to P4S methods using prior information like 2D drawings and 3D models [29]. These methods have been applied to civil infrastructures [50], space frame structures [29], rebar detection [27, 28], and landslide monitoring [54]. In cases where prior information is outdated or unavailable, low-resolution preliminary scans are required, either indoors [39] or outdoors using UAVs [26].
P4S optimizations aim to maximize coverage and overlap while minimizing the number of viewpoints [19, 7]. Methods like CMA Evolution Strategy [40] and regular grid sampling [21, 33] increase the likelihood of finding optimal viewpoints but also raise computational complexity. Jia’s approach using grid-based sampling [20, 21, 22] scores candidate viewpoints based on observed wall segments, though this approach is limited by its reliance on reference objects for registration. Noichl [35] improve Jia’s method into 3D scene for an uniform covering by triangulating 3D scene into meshes and checking their visibility. For scenarios lacking BIM priors, structural elements can be surveyed and detected using low-cost sensors to serve as a preliminary prior for static station planning. A precise scan is then performed in a stop-and-go manner using a static scanner mounted on a robotic dog [9, 18].
Our approach addresses these limitations by using a continuous Visibility Field (VF) model, which reduces optimization from 2D to 1D by focusing on structural convergence points like medial axis and joints, thus minimizing computation and enhancing adaptability.
2.2 Mobile-Based Exploration
Coverage Path Planning (CPP) focuses on computing optimal, collision-free paths for robots to fully cover target areas [44]. Techniques like Rapidly-exploring Random Trees (RRT) and their variants (RRT* [23], RRT# [37]) are widely used for autonomous exploration in unknown spaces, including UAVs [31] and wheeled robots [38]. RRT-based methods maximize information gain while minimizing travel distance [17, 47, 53].
Methods utilizing tensor fields for guiding robot movements enable efficient path routing in partially reconstructed indoor scenes [48, 46]. Recent approaches like the Signed Distance Field (SDF) and Hamilton-Jacobi skeleton further improve path planning in known environments [36]. These concepts share similarities with our visibility field, which leverages converging lines for efficient viewpoint planning.
For known environments, algorithms like the receding horizon ”next-best-view” approach [3] and dual-resolution mapping schemes [6] balance detailed local mapping with efficient global exploration. For applications in autonomous driving and exploration, recent advances such as ”LookOut” [12] propose diverse multi-future prediction models that enhance autonomous navigation by predicting possible trajectories, balancing safety, and efficiency.
3 Theoretical Foundations and Global Optimality
This section formulates the Viewpoint Planning Problem (VPP) for static LiDAR scanning in a 2D polygonal environment and establishes the theoretical basis for restricting candidate viewpoints to the medial axis (MA). We first present the problem definition, including coverage and connectivity requirements, and then show that any line segment visible from an interior point is also visible from a corresponding set of MA points. This demonstrates that restricting the viewpoint search space to the MA incurs no loss of coverage capability. Finally, we prove that a minimal solution confined to the MA is globally optimal.
3.1 Problem Definition
Let $P$ be a 2D polygonal region (such as a floor plan) with possible obstacles or holes, while remaining topologically connected. The polygon’s boundary (including internal obstacles) is discretized into $n$ small line segments:
$L=\{\,l_{1},\;l_{2},\;\dots,\;l_{n}\},$
which represent the target structures (e.g., walls or windows) to be covered by static LiDAR scanners (hereafter referred to as viewpoints).
Candidate Viewpoints
We aim to position a set of viewpoints such that all segments in $L$ are fully visible, while the chosen viewpoints remain connected under an overlap criterion. Let $V=\{\,v_{1},\;v_{2},\;\dots,\;v_{m}\}$ be the set of candidate viewpoint locations. We will later establish that it is sufficient to confine $V$ to the medial axis of $P$, reducing a 2D search to a simpler 1D structure.
Visibility
we define a Completeness Table $C$, where $C_{ij}$ indicates whether viewpoint $V_{i}$ observes line segment $L_{j}$
$C_{ij}=\begin{cases}1&\text{if }v_{i}\text{ has an unobstructed view of }l_{j}\\ 0&\text{otherwise}\end{cases}$ (1)
Connectivity via Overlap
Consider any pair of viewpoints $(v_{i},v_{k})$. Let $O_{ik}$ be their overlap ratio, i.e., the fraction of segments in $L$ that both $v_{i}$ and $v_{k}$ can simultaneously observe. For a threshold $\tau$, an edge $(v_{i},v_{k})$ exists in an overlap graph $G=(V,E)$ if $O_{ik}\geq\tau$. A chosen subset $S\subseteq V$ must then form a connected subgraph under this threshold.
Mathematical Formulation
We seek a minimal subset $S\subseteq V$ of viewpoints such that:
  1. (a)
    Coverage: Each segment $l_{j}\in L$ is observed by at least one viewpoint in $S$, i.e.,
    $\forall\,l_{j}\in L,\;\exists\,v_{i}\in S:\;C_{ij}=1.$ (2)
  2. (b)
    Connectivity: The induced subgraph $(S,E_{S})$ is connected under the overlap threshold $\tau$, where
    $E_{S}\;=\;\bigl{\{}(v_{i},v_{k})\in E\;\big{|}\;v_{i},v_{k}\in S,\;O_{ik}\geq\tau\bigl{\}}.$ (3)
The optimization problem becomes:
[…]
Since problem (4) generalizes the set cover problem (with additional connectivity constraints), it remains NP-hard. Nonetheless, by restricting $V$ to the medial axis (MA), we retain optimality while reducing the search complexity considerably.
3.2 Skeleton Completeness and Visibility Propagation
Medial Axis (MA)
The MA of $P$ is the set of points inside $P$ that are equidistant to at least two boundary edges. In Blum’s grassfire transform [4], if the boundary of $P$ is ignited simultaneously, the inward-propagating flames collide and form the locus of the MA.
Theorem 1 (Skeleton Completeness).
Suppose an interior point $p\in P$ sees a set of boundary segments $L_{p}\subseteq L$. Then there exists a set of points $\{q_{l}\}\subseteq\mathrm{MA}(P)$ such that
$L_{p}\;\;\subseteq\;\;\bigcup_{l\in L_{p}}\mathrm{Vis}(q_{l}).$
Hence, for any interior coverage provided by $p$, there exists an MA-based set of points whose collective coverage reproduces that of $p$.
Sketch of Proof. Using Blum’s “fire transformation,” the visibility from every segment propagates inward until it intersects the medial axis $\mathrm{MA}(P)$. Each line segment $l_{j}\in L_{p}$ must intersect $\mathrm{MA}(P)$ at some point $q_{j}$ visible to $l_{j}$. Thus,
$\mathrm{Vis}(p)\;\;\subseteq\;\;\bigcup_{l_{j}\in\mathrm{Vis}(p)}\,\mathrm{Vis}(q_{j}).$
Consequently, although a single point on the MA may not match $p$’s entire coverage, a suitable collection of MA points can, illustrated in 2. Restricting $V$ to $\mathrm{MA}(P)$ therefore does not diminish the achievable coverage.
Figure 2: Illustration of skeleton completeness. (a) A simple polygon (black boundary) with its medial axis (gray arcs). (b) Coverage by an interior viewpoint $p$ can be replicated by a single skeletal viewpoint $q$ lying on the medial axis (MA). (c) In the general case, coverage from an interior viewpoint $p$ may require multiple skeletal viewpoints (here, $m$ and $n$) to fully reproduce $p$’s visibility. Thus, no coverage capability is lost by restricting candidate viewpoints to the MA. For clarity, boundary segments commonly observed by both $p$ and the MA viewpoints ($q$, or $m$ and $n$) are not drawn.
3.3 Global Minimal Cover–Connectivity on the Skeleton
By Theorem 1, any coverage by interior points can be equivalently transferred to a suitable set of points on the medial axis. Hence, in the formulation (4), restricting $V$ to $\mathrm{MA}(P)$ (including any branch or joint nodes) captures all feasible coverage solutions.
Theorem 2 (Global Minimal Cover–Connectivity).
Let $S^{*}\subseteq\mathrm{MA}(P)$ be a minimal solution to (4) that covers all $l_{j}\in L$ and satisfies the overlap-based connectivity requirement. Then $S^{*}$ is irreducible (i.e., removing any $v_{i}\in S^{*}$ breaks coverage or connectivity), and no smaller interior configuration can outperform $S^{*}$. Therefore, $S^{*}$ is globally optimal.
Implication and Heuristic Search
Exploiting the symmetry of the medial axis, no point outside $\mathrm{MA}(P)$ can provide superior coverage or connectivity. Consequently, a minimal cover–connectivity solution restricted to the skeleton cannot be improved by adding non-skeleton points. Hence, solving (4) directly on $\mathrm{MA}(P)$ yields a globally optimal arrangement in $P$. Although the problem remains NP-hard, standard polynomial-time greedy algorithm (e.g., [10]) often produce near-optimal results in practice. Once obtained, the minimal MA-based solution is irreducible, further affirming its global optimality for the entire polygon.
Figure 3: Overview of the Viewpoint Network (VPN) optimization based on the Visibility Field (VF). (a) Input floor plan; (b) Visibility Field (VF); (c) Distance field of VF; (d) Converging lines and points; (e) Optimized Viewpoint Network.
4 Method
4.1 Overview
Our proposed VF-Plan addresses the NP-hard complexity of the VPP through a greedy approach centered on the VF. Given a structural floor plan, VF-Plan constructs a Viewpoint Network (VPN) that achieves comprehensive coverage with minimal viewpoints, ensuring connectivity and registrability. Each viewpoint is selected to form a fully connected observation network, minimizing registration errors via Least Squares Fitting (LSF).
As illustrated in Figure 3, the VF-Plan pipeline starts from a floor plan with structural elements (walls, columns, doors, windows) and proceeds through the steps of VF computation, distance field generation, identification of medial axis and joint points, and initialization of candidate viewpoints. A greedy algorithm then refines these candidates into a minimal, connected VPN. To optimize the computationally intensive task of determining visible wall segments from each viewpoint, we employ a BSP-tree[14] to accelerate processing.
4.2 Visibility Field
Laser scanning is limited to visible surfaces, leaving occluded areas unobserved. Therefore, visibility and coverage analysis are essential for viewpoint planning, with key objectives of maximizing observation completeness and registrability. The field of view is central to VPN optimization, defining the information captured at each viewpoint and guiding the selection of optimal viewpoints.
A common method for evaluating the information scanned by a LiDAR sensor is based on the length of observed walls [22]. Typically, wall length is counted either by ensuring both endpoints are visible or by partitioning walls into uniformly sampled segments within the scanner’s range. For 3D cases, visibility is evaluated by counting the number of observed voxels. However, these approaches lack normalization, yielding values that vary with discretization resolution and do not account for differences in distance or orientation to the scanner, impacting computational stability and load without improving coverage accuracy.
Our approach instead uses the valid scanned angle to integrate relevant factors into a single index, uniformly scaled from 0 to $2\pi$. We define a static LiDAR scanner with a 360° field of view in the X-Y plane, a range from $r_{\text{min}}$ to $r_{\text{max}}$, and uniform scanning speed, as shown in Figure 4. Only objects within this range are observable. Generally, closer viewpoints collect more information, but excessive proximity can occlude other parts of the field. Scanning is performed in discrete observation angles, with observed information defined as the valid observed angle, ensuring alignment with scanner properties.
Figure 4: Scan model of a static scanner. The scanner provides valid observations within a 360° angle and a range from $r_{\text{min}}$ to $r_{\text{max}}$.
To quantify observed data, let Circle_min intersect line segment $AB$ at points $C$ and $D$, and Circle_max at points $E$ and $F$. The valid observed line segment is:
$L_{\text{valid}}=\cap(L_{AB},L_{EF})-\cap(L_{AB},L_{CD})$ (5)
and the corresponding observed angle is:
$\theta_{\text{valid}}=\text{Angle}_{O}(L_{\text{valid}})$ (6)
Each viewpoint is assigned a View Angle, with the scene represented as a grid where each cell center is a viewpoint. Thus, the scene is represented as a Visibility Field, as shown in Figure 3(b), which aligns well with LiDAR properties and expert experience, displaying expected visibility patterns.
4.3 Converging Lines and Joints
Observation data naturally converge to the medial axis and then joints of the Visibility Field, so placing viewpoints along these skeletons and joints ensures full coverage. Thus, VPN optimization reduces from a 2D to a 1D problem along these skeletons. We use an medial axis algorithm [43] to identify skeletons and joint points, placing candidate viewpoints along these lines. Where overlap ratios are insufficient, additional viewpoints are added to ensure network cohesion.
For any skeleton ridge $R_{ij}$ connecting joints $J_{i}$ and $J_{j}$, we calculate an Overlap Ratio $O_{ij}$ of the endpoints. When $O_{ij}$ falls below a threshold, a midpoint $J_{m}$ is added to split the ridge, ensuring connectivity. Both the skeleton joints and connecting points are called as converging points. And the resulting skeleton ridges are called converging lines. This iterative placement process continues until all converging points have adequate overlap for registrability.
4.4 Overlap Ratio
Figure 5: Overlap calculation between observed segments from different viewpoints, quantifying connectivity.
The Overlap Ratio quantifies shared visibility between viewpoints, a requirement for robust registration. Five overlap ratios based on length and angle are defined, ensuring effective connectivity independent of distance alone.
Given wall segments observed by viewpoint $V_{a}$ as $L_{a}$ with an observation angle $\theta_{a}$:
$\theta_{a}=\text{Angle}(V_{a},L_{a})$ (7)
The segments observed by both $V_{a}$ and $V_{b}$ are $L_{ab}$:
$L_{ab}=L_{a}\cap L_{b}$ (8)
The length-based overlap ratios are:
$O^{\text{Min\_Len}}=\frac{L_{ab}}{\min(L_{a},L_{b})}$ (9)
$O^{\text{Mean\_Len}}=\frac{2L_{ab}}{L_{a}+L_{b}}$ (10)
$O^{\text{Union\_Len}}=\frac{L_{ab}}{L_{a}\cup L_{b}}$ (11)
And the angle-based overlap ratios are:
$O^{\text{Union\_Ang}}=\frac{\theta_{a}^{ab}+\theta_{b}^{ab}}{\theta_{a}+\theta_{b}}$ (12)
$O^{\text{Mean\_Ang}}=\frac{\theta_{a}^{ab}}{2\theta_{a}}+\frac{\theta_{b}^{ab}}{2\theta_{b}}$ (13)
4.5 Greedy Optimization for VPN Construction
The VF-Plan greedy algorithm constructs an efficient Viewpoint Network (VPN) by starting with candidate viewpoints positioned strategically at key joints within the Visibility Field. For each pair of candidate viewpoints, an Overlap Ratio $O$ is calculated, with pairs showing adequate overlap retained as adjacencies in the VPN to ensure robust connectivity.
Figure 6: Greedy Optimization of VPN. (a) Floorplan showing candidate viewpoints (labeled a to k) positioned along walls (numbered 0 to 17); (b) Connectivity graph, where edges link viewpoints with overlap ratio $O$ exceeding the threshold criterion; (c) Initial seed viewpoint selection based on maximum wall visibility (numbers near viewpoints indicate the count of observed walls); (d) - (f) Iterative expansion of the seed viewpoint set by adding adjacent viewpoints that provide the highest additional coverage, enhancing overall viewpoint arrangement.
To systematically track visibility, we construct the Completeness Table $C$ (see Equation 1), which determines whether a given viewpoint $v_{i}$ observes a specific line segment $l_{j}$. To enhance precision in observation tracking, wall boundaries are uniformly partitioned into equal-length segments.
Algorithm 1 Greedy Optimization of VPN
Input: Candidate viewpoints $V$, line segments $L$
Output: Minimal set of viewpoints $V_{\text{opt}}$ covering all $L$ and forming a connected network
  1. 1.
    Initialize Completeness Table $C$
  2. 2.
    Identify $V_{i}$ with maximum $C_{i}$, add $V_{i}$ to Viewpoint Seed Set (VPS)
  3. 3.
    Add adjacent viewpoints of $V_{i}$ to VPB
  4. 4.
    Remove all $L$ observed by $V_{i}$ from $C$
  5. 5.
    While unobserved $L$ exists in $C$:
    1. (a)
      Select $V_{i}$ in VPB maximizing new coverage $C_{i}$
    2. (b)
      If multiple options, select $V_{i}$ with highest overlap $O$ with VPS
    3. (c)
      Move $V_{i}$ to VPS, add adjacencies to VPB
    4. (d)
      Remove all $L$ observed by $V_{i}$ from $C$
return VPS as the optimized minimal VPN
As outlined in Algorithm 1 and illustrated in Figure 6, the greedy optimization initiates by selecting the viewpoint with the highest coverage as the seed. The algorithm then iteratively expands by adding neighboring viewpoints that maximize additional coverage, ensuring each new viewpoint significantly contributes to covering line segments while constructing a minimal, fully connected Viewpoint Network (VPN).
Upon completion of the initial search, the algorithm produces a single connected component encompassing all essential viewpoints, with sufficient overlap to facilitate stable connections. The algorithm subsequently enhances network stability by identifying cycles within the VPN graph. For nodes that cannot form cycles (i.e., loose nodes), Converging Points are added to ensure cohesion without redundancy. Additionally, any node with more than three connecting edges from a cycle selects a Converging Joint to establish loops, thereby strengthening network connectivity.
5 Experiments
Figure 7: Performance of five proposed overlap ratios. The first row shows overlap ratios between moving and reference viewpoints; the second row illustrates room configurations, with black arrows indicating movement direction and green dots marking reference viewpoints.
5.1 Dataset
In this section, we evaluate the proposed VPN optimization algorithm across multiple scenarios and compare its performance with state-of-the-art methods in both indoor and outdoor environments.
Due to the lack of a dedicated benchmark for VPN optimization, we compiled a dataset from established sources, focusing on two primary scene types: indoor and outdoor environments. This dataset provides a robust foundation for evaluating VPN optimization across a range of realistic settings.
For indoor environments, we selected scenes with varied room layouts to evaluate the algorithm’s adaptability across multiple configurations. The Structure3D dataset [56] was included for its extensive collection of diverse Asian residential scenes, featuring both complex and compact room types frequently used in 3D modeling and synthetic data generation. VPNs generated from this dataset can also facilitate synthetic LiDAR data production for advancing research in 3D scene understanding. Specific scenes such as 00007, 00009, 00071, 02970, and 02972, discussed in subsequent sections, are drawn from Structure3D. For further comparisons, we incorporated additional scenes from Zeng1 [51], Noichl [34], Xu [48], and TUB1 from the ISPRS benchmark [52, 24].
For outdoor scenarios, we limited our approach to 2D site plans without considering terrain elevation, distinguishing our method from elevation-based approaches like [8]. Selected scenes include the CCIT and Crowsnest datasets from the University of Calgary [20] and the ECUT dataset from East China University of Technology [8], which represent open-space environments suitable for VPN optimization.
To evaluate VPN quality, we rely on expert-labeled annotations as standard ground truth data is unavailable. Where possible, we use existing expert-labeled VPNs (e.g., Zeng1, scene1 and TUB1), and for other cases, independent expert labels were created for comparison with manual designs.
5.2 Performance Metrics
The optimization algorithm aims to construct a minimal, fully connected viewpoint network with reduced redundancy. As most state-of-the-art methods achieve 100% coverage, we focus on two key metrics: the number of optimal viewpoints and network connectivity.
Network connectivity is measured using the Weighted Average Path Length (WAPL), which reflects the compactness of the network [5]. A shorter WAPL indicates stronger connectivity and a more efficient structure. For a network with weighted edges $w$, WAPL is calculated as:
$WAPL=\frac{1}{N(N-1)}\sum_{i\neq j}d_{w}(i,j)$ (14)
where $d_{w}(i,j)$ represents the shortest weighted path length between nodes $i$ and $j$, and $N$ is the total number of nodes.
The edge weight $w$ is defined as
$w_{ij}=1-O_{ij}$ (15)
A larger $O_{ij}$ implies stronger registration-based connectivity, meaning a shorter length between the two nodes $i$ and $j$.
In cases where baseline methods do not yield a single connected component, some viewpoints may have a small overlap ratio with others. To ensure WAPL remains computable, we assign a large weight (100) to node pairs that lack a connecting path.
5.3 Algorithm Parameter Effects
5.3.1 Overlap Ratio
We propose five overlap ratios according to the length of scanned walls and according angles to viewpoints in section 4.4. Here we evaluate them in two test indoor scenes, illustrated in Figure 7. Among them, $O^{\text{Min\_Len}}$ quickly reaches 100%, showing limited variability along movement paths, as seen from B to H in Figures 7(b) and (e), making it less suitable for dynamic scenes. The angle-based measures, $O^{\text{Union\_Ang}}$ and $O^{\text{Mean\_Ang}}$, exhibit non-monotonic behavior, particularly along paths from C to E, shown in Figures 7(c) and (f), which limits their usefulness for consistent registration. The $O^{\text{Union\_Len}}$ measure tends to be too small, often underestimating connectivity. Therefore, we select $O^{\text{Mean\_Len}}$ for subsequent experiments and analyses due to its balanced performance across different scenarios.
5.3.2 Scan Radius
The scan model used in this study operates in the X-Y plane with a 360° field of view, bounded by $R_{\text{min}}$ and $R_{\text{max}}$, defining near and far blind zones. Typical static scanners, including the FARO M70, Trimble X7, and Leica BLK360, have maximum ranges around 40 meters and minimum ranges of about 1 meter, allowing for comprehensive indoor coverage but requiring precise placement in small rooms and narrow corridors.
For outdoor scenarios, we examine the impact of varying $R_{\text{max}}$, testing values at 15m, 30m, 45m, and 75m, with $R_{\text{min}}$ fixed at 1m. As illustrated in Figure 8, increasing $R_{\text{max}}$ expands the area of valid placements for viewpoints—defined where the View Angle exceeds 1.0—and reduces the Viewpoint Count (VC) needed for optimized coverage, from 23 to 14, 10, and 8. Larger $R_{\text{max}}$ values are advantageous for open spaces, while smaller values require additional viewpoints for full coverage. These results indicate that our method aligns effectively with real-world scanner specifications and expert guidelines.
Figure 8: Effect of $R_{\text{max}}$ on Viewpoint Placement and Network Efficiency. As $R_{\text{max}}$ increases from 15m to 75m, the valid placement area (where View Angle $>$ 1.0) grows, reducing the Viewpoint Count (VC) needed for optimized VPN coverage from 23 to 8.
5.3.3 Wall Partition
To improve station coverage accuracy, walls are partitioned into segments of varying lengths, as shown in Figure 9. We find that partitioning does not alter the final optimization result, but only affects the selection order of stations. Larger partition lengths (or no partitioning) place initial viewpoints at major intersections, optimizing computational cost and maintaining effective coverage.
Figure 9: Effects of wall partitions. From left to right: unpartitioned walls and walls partitioned into segments of 1.0m, 0.5m, 0.1m, and 0.01m.
5.3.4 Effect of Windows
Including windows in the model adds complexity to the skeleton structure, generating more detailed skeletons with additional joint points, as shown in Figure 10. When windows are considered, the VPN algorithm requires only a minimal increase in viewpoints to achieve complete coverage, as seen in Table 1.
Without windows, the Viewpoint Count (VC) averages around 1.4 times the number of rooms, while with windows, it increases slightly to 1.6 times the room count. These results indicate that our algorithm effectively accommodates the added detail from windows. For precise window scans, including window frames in the optimization process is recommended.
Figure 10: Comparison of VPN structures with and without windows. Including windows adds complexity to the skeleton and slightly increases the Viewpoint Count.
Scene 00007 00009 00071
Number of Rooms 5 3 9
VC (Without Windows) 7 4 14
VC (With Windows) 7 5 15
Table 1: Viewpoint Count (VC) comparison for room coverage with and without windows across different scenes.
5.4 Results and Analysis
Scene WAPL VC Baseline Method
Expert Ours Baseline Expert Ours Baseline
Zeng1 0.47 0.53 43.37 12 12 10 Zeng[51]
Noichl 0.32 0.38 23.91 8 10 7 Noichl2023[34]
TUB1 1.03 1.00 40.97 28 29 19 Zhai2024[52]
Xu 0.35 0.31 28.69 8 8 7 Xu2017[48]
CCIT 0.52 0.47 38.12 8 8 7 Jia2018[20]
Crowsnest 0.50 0.49 34.02 9 8 8 Jia2018[20]
ECUT 0.67 0.76 11.94 58 52 139 Chen2023[8]
Table 2: Comparison of viewpoint counts (VC) and Weighted Average Path Length (WAPL) across indoor and outdoor scenes. Lower WAPL indicates more compact network connectivity.
We benchmarked our method against state-of-the-art approaches using datasets from prior studies, representing a range of indoor and outdoor scenes. Indoor datasets include Zeng1 [51], Noichl [34], and TUB1 from the ISPRS dataset [52]. Outdoor datasets consist of CCIT and Crowsnest from the University of Calgary [20] and ECUT from East China University of Technology [8].
Indoor Results. As shown in Figure 12, our method’s viewpoint optimization closely aligns with expert-designed networks, achieving comprehensive coverage and creating compact, interconnected structures. This optimized network design enhances registration accuracy and connectivity, even in complex indoor settings.
Outdoor Results. For outdoor scenes, our method effectively adapts to open spaces, as illustrated in Figure 13. Adjustments to the $R_{\text{max}}$ parameter notably improved performance in expansive environments, reducing the required number of viewpoints without compromising coverage.
Table 2 demonstrates that our method performs comparably or better than expert-designed networks, achieving similar viewpoint counts (VC) and lower Weighted Average Path Length (WAPL). This outcome indicates a more compact and cohesive network structure. Unlike most baselines that often yield fragmented networks (highlighted by green ellipses in Figures 12 and 13), our approach maintains low WAPL scores, which reflect enhanced connectivity and minimized redundancy.
In indoor environments, our method achieves a 98% reduction in WAPL compared to baseline methods, creating highly connected networks well-suited for confined spaces. For complex outdoor scenarios, such as ECUT, our approach achieves a 91% WAPL reduction over baselines, demonstrating scalability and adaptability across both intricate indoor layouts and expansive outdoor scenes.
To further validate the optimized viewpoints, we simulated static LiDAR scanning at each optimized viewpoint using Helios++ [13]. As shown in Figure 11, the simulated LiDAR point clouds confirm complete scene coverage, with effective overlap between adjacent scans across both indoor and outdoor environments.
Figure 11: Simulated LiDAR points from optimal viewpoints. Each scan is uniquely colored. Top row shows top views, and bottom row shows oblique views with tripods indicating viewpoints.
Figure 12: Viewpoint planning results for indoor environments. Red dots represent individual viewpoints, while green ellipses indicate clusters of viewpoints that are not connected to the main network.
Figure 13: Viewpoint planning results for outdoor environments. Red dots represent individual viewpoints, while green ellipses indicate clusters of viewpoints that are not connected to the main network.
6 Conclusions
We introduced a novel solution to the Viewpoint Planning Problem and Art Gallery Problem in static LiDAR scanning by leveraging a Visibility Field to reduce optimization complexity from 2D to 1D. This VF-based approach enables the creation of a minimal, fully connected viewpoint network that ensures comprehensive coverage with reduced redundancy, advancing viewpoint planning for static LiDAR applications.
Our method offers a scalable and efficient framework applicable in structured and semi-structured environments, addressing real-world challenges in 3D data collection and autonomous navigation. Our VF framework adapts well to real-world settings with potential obstructions by integrating partial occlusion handling based on initial scans.
In summary, this method provides a practical and robust solution for static viewpoint planning, with promising extensions to dynamic environments. Future work will focus on its application for large-scale LiDAR data collection and synthetic data generation, and expanding the method to support path planning in both 2D and 3D scenarios for more versatile deployment in complex environments.
CRediT authorship contribution statement
Biao Xiong: Writing - Original draft, Validation, Methodology, Investigation, Project administration. Longjun Zhang: Original draft, Validation, Software, Methodology. Ruiqi Huang: Validation, Software, Methodology, Data curation. Junwei Zhou: Validation, Software, review & editing. Bojian Wu: Writing - review & editing, Writing - original draft, Visualization. Fashuai Li: Writing - review & editing, Supervision, Resources, Conceptualization.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@
 Abstract
-Viewpoint planning is crucial for 3D data collection and autonomous navigation, yet existing methods often miss key optimization objectives for static LiDAR, resulting in suboptimal network designs.
-The Viewpoint Planning Problem (VPP), which builds upon the Art Gallery Problem (AGP), requires not only full coverage but also robust registrability and connectivity under limited sensor views.
-We introduce a greedy optimization algorithm that tackles these VPP and AGP challenges through a novel Visibility Field (VF) approach.
-The VF captures visibility characteristics unique to static LiDAR, enabling a reduction from 2D to 1D by focusing on medial axis and joints.
-This leads to a minimal, fully connected viewpoint network with comprehensive coverage and minimal redundancy.
-Experiments across diverse environments show that our method achieves high efficiency and scalability, matching or surpassing expert designs.
-Compared to state-of-the-art methods, our approach achieves comparable viewpoint counts (VC) while reducing Weighted Average Path Length (WAPL) by approximately 95%, indicating a much more compact and connected network.
-Dataset and source code will be released upon acceptance.
-keywords:
-Viewpoint Optimization , Terrestrial Laser Scaning , Visibility Field , Medial axis transform , Greedy searching , BIM
+Viewpoint planning is critical for efficient 3D data acquisition in applications such as 3D reconstruction, building life-cycle management, navigation, and interior decoration.
+However, existing methods often neglect key optimization objectives specific to static LiDAR systems, resulting in redundant or disconnected viewpoint networks.
+The viewpoint planning problem (VPP) extends the classical Art Gallery Problem (AGP) by requiring full coverage, strong registrability, and coherent network connectivity under constrained sensor capabilities.
+To address these challenges, we introduce a novel Visibility Field (VF) that accurately captures the directional and range-dependent visibility properties of static LiDAR scanners.
+We further observe that visibility information naturally converges onto a 1D skeleton embedded in the 2D space, enabling significant searching space reduction.
+Leveraging these insights, we develop a greedy optimization algorithm tailored to the VPP, which constructs a minimal yet fully connected Viewpoint Network (VPN) with low redundancy.
+Experimental evaluations across diverse indoor and outdoor scenarios confirm the scalability and robustness of our method.
+Compared to expert-designed VPNs and existing state-of-the-art approaches, our algorithm achieves comparable or fewer viewpoints while significantly enhancing connectivity.
+In particular, it reduces the weighted average path length by approximately 95%, demonstrating substantial improvements in compactness and structural efficiency.
+Code is available at https://github.com/xiongbiaostar/VFPlan.
+{CCSXML}
+<ccs2012> <concept> <concept_id>10003752.10010061.10010063</concept_id> <concept_desc>Theory of computation Computational geometry</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
 […]
-[label1]organization=School of Computer Science and Artificial Intelligence, Wuhan University of Technology,addressline=122 Luoshi Rd, city=Wuhan, postcode=430070, state=Hubei, country=China
+[500]Theory of computation Computational geometry
 […]
-[label2]organization=College of Computer Science and Technology, Zhejiang University, addressline=866 Yuhangtang Rd, city=Hangzhou, postcode=310058, state=Zhejiang, country=China
-[…]
-[label3]organization=The Advanced Laser Technology Laboratory of Anhui Province, addressline= 850 Huangshan Rd, city=Hefei, postcode=230031, state=Anhui, country=China
-{graphicalabstract}
-Viewpoint optimization for static LiDAR Scanning.
-(a) The input floor plan; (b) Optimized viewpoints indicated by green dots, with the Visibility Field displayed as the background image; (c) and (d) Simulated LiDAR points from optimal viewpoints, shown in 2D and 3D respectively, with each scan distinguished by a unique color.
-{highlights}
-Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the Art Gallery Problem, achieving efficient coverage with minimal viewpoints and robust network connectivity.
-Visibility Field for Dimensionality Reduction: We introduce a continuous Visibility Field tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
-Greedy Algorithm for Efficient Coverage: Our greedy Algorithm utilizes the VF’s structural insights to construct a minimal, fully connected viewpoint network, ensuring both robust coverage and network connectivity.
 1 Introduction
-The Viewpoint Planning Problem (VPP) has emerged as a critical research spot due to the growing use of LiDAR and camera sensors in applications such as 3D reconstruction, autonomous navigation, and Scan-to-BIM applications [41, 30].
-VPP focuses on determining the optimal placements and configurations of sensors to achieve specific reconstruction goals, such as maximizing observation completeness and minimizing occlusions, while balancing efficiency and data quality.
-Both model-free and model-based algorithms for viewpoint and path planning have gained significant attention due to their applicability across various industries, including robotics, construction, and surveillance [31].
-Conceptually, VPP is rooted in the Art Gallery Problem (AGP), a classic challenge in computational geometry that seeks the minimal number of “guards” required to cover a given area, typically a polygon, with full 360° visibility [25].
+The VPP has emerged as a critical research spot due to the growing use of LiDAR and camera sensors in applications such as 3D reconstruction, autonomous navigation, facility assessment, building life-cycle management, and interior decoration applications [SRR03, MLL23].
+VPP focuses on determining the optimal placements and configurations of sensors to achieve specific reconstruction goals, such as maximizing observation coverage and minimizing occlusions, while balancing efficiency and data quality.
+Both model-free and model-based algorithms for viewpoint and path planning have gained significant attention due to their applicability across various industries, including robotics, construction, and surveillance [MHS23].
+Conceptually, VPP is rooted in the AGP, a classic challenge in computational geometry that seeks the minimal number of “guards” required to cover a given area, typically a polygon, with full 360° visibility [LL86].
 AGP is often simplified by constraining guard placements to polygon vertices; however, this simplification does not capture the true computational complexity of the problem.
-In its generalized form, AGP is $\exists\mathbb{R}$-complete, meaning it is at least as hard as any NP-complete problem, with no known efficient solutions [11, 1].
-This NP-hard nature has inspired substantial research interest, particularly in constrained versions that make idealized assumptions, such as complete visibility and static scenes [11, 41].
-Figure 1: Viewpoint optimization for static LiDAR Scanning.
-(a) The input floor plan; (b) Optimized viewpoints indicated by green dots, with the Visibility Field displayed as the background image; (c) and (d) Simulated LiDAR points from optimal viewpoints, shown in 2D and 3D respectively, with each scan distinguished by a unique color.
+In its generalized form, AGP is $\exists\mathbb{R}$-complete, meaning it is at least as hard as any NP-complete problem, with no known efficient solutions [CdRdS11, AAM21].
+This NP-hard nature has inspired substantial research interest, particularly in constrained versions that make idealized assumptions, such as complete visibility and static scenes [CdRdS11, SRR03].
 VPP, however, introduces additional complexities beyond those of AGP.
 Unlike static guards, sensors in VPP must maintain overlapping fields of view for accurate data registration, accommodate limited and directional visibility ranges, and meet high accuracy requirements for 3D scene reconstruction.
-These real-world demands, especially in cluttered or partially obstructed environments, render VPP a more challenging NP-hard problem, requiring innovative optimization strategies beyond those used for AGP [41, 1].
-Our work addresses these challenges by introducing a novel greedy approach based on the concept of a Visibility Field (VF), specifically designed for static LiDAR applications.
-The VF captures the unique visibility characteristics of LiDAR, including constraints on range and incident angles, enabling a significant reduction in computational complexity.
-By focusing on critical points such as medial axis and joints within the VF, we effectively reduce the optimization space from 2D to 1D, facilitating the construction of a minimal, fully connected viewpoint network that achieves comprehensive coverage with reduced redundancy.
-Our main contributions are as follows:
+These real-world demands, especially in cluttered or partially obstructed environments, render VPP a more challenging NP-hard problem, requiring innovative optimization strategies beyond those used for AGP [SRR03, AAM21].
+We address the challenges of static LiDAR coverage planning by introducing a novel greedy algorithm based on the concept of a Visibility Field (VF).
+The VF encapsulates the physical constraints of LiDAR sensors, including limited sensing range and angle-dependent visibility, enabling a substantial reduction in computational complexity.
+We formally demonstrate that visibility information is inherently concentrated along a one-dimensional skeleton embedded within the two-dimensional space.
+This observation allows us to restrict the search space from 2D to 1D by targeting critical structural features such as the medial axis and junctions within the VF.
+Consequently, we construct a minimal, fully connected network of viewpoints that achieves complete scene coverage with reduced redundancy.
+Our main contributions are summarized as follows:
   1. 1.
-    Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the Art Gallery Problem, achieving efficient coverage with minimal viewpoints and robust network connectivity.
+    Unified Solution to VPP and AGP: Our method addresses the combined challenges of viewpoint planning and the AGP, achieving efficient coverage with minimal viewpoints and robust network connectivity.
   2. 2.
-    Visibility Field for Dimensionality Reduction: We introduce a continuous Visibility Field tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
+    Visibility Field for Dimensionality Reduction: We introduce a continuousVF tailored to static LiDAR systems, reducing the optimization space from 2D to 1D by leveraging structural elements like medial axis and joints.
   3. 3.
-    Greedy Algorithm for Efficient Coverage: Our greedy algorithm utilizes the VF’s structural insights to construct a minimal, fully connected viewpoint network, ensuring both robust coverage and network connectivity.
+    Greedy Algorithm for Efficient Coverage: Our greedy algorithm utilizes the VF’s structural insights to construct a minimal, fully connected VPN, ensuring both robust coverage and network connectivity.
+Figure 1: Pipeline of viewpoint network (VPN) optimization using VF.
+(a) Input floor plan with walls, doors, and windows; (b) Computed VF encoding the valid observed angle at each pixel; (c) Distance field derived from the VF; (d) Extraction of converging lines (medial axis) and converging points (joints) used as candidate viewpoints; (e) Final optimized VPN ensuring full coverage and strong connectivity.
 2 Related Work
@@
 2.1 Station-Based Scanning
-Sensor network design shares similarities with the Planning for Scanning (P4S) problem, observed in applications like surveillance cameras [32, 55, 16], directional sensors [15], 5G base stations [42, 45], stealth game [49], and station-based LiDAR scanning [2].
-For LiDAR, the goal is to cover the entire scene with comprehensive wall scanning while ensuring data quality criteria such as completeness and accuracy [29, 2].
-Advancements in LiDAR technology have led to P4S methods using prior information like 2D drawings and 3D models [29].
-These methods have been applied to civil infrastructures [50], space frame structures [29], rebar detection [27, 28], and landslide monitoring [54].
-In cases where prior information is outdated or unavailable, low-resolution preliminary scans are required, either indoors [39] or outdoors using UAVs [26].
-P4S optimizations aim to maximize coverage and overlap while minimizing the number of viewpoints [19, 7].
-Methods like CMA Evolution Strategy [40] and regular grid sampling [21, 33] increase the likelihood of finding optimal viewpoints but also raise computational complexity.
-Jia’s approach using grid-based sampling [20, 21, 22] scores candidate viewpoints based on observed wall segments, though this approach is limited by its reliance on reference objects for registration.
-Noichl [35] improve Jia’s method into 3D scene for an uniform covering by triangulating 3D scene into meshes and checking their visibility.
+Sensor network design shares similarities with the Planning for Scanning (P4S) problem, observed in applications like surveillance cameras [MADR11, ZYCH13, GCW15], directional sensors [FG09], 5G base stations [SSZ20, WZL20], stealth game [XTV14], and station-based LiDAR scanning [ABT21].
+For LiDAR, the goal is to cover the entire scene with comprehensive wall scanning while ensuring data quality criteria such as completeness and accuracy [LXHL21, ABT21].
+Advancements in LiDAR technology have led to P4S methods using prior information like 2D drawings and 3D models [LXHL21].
+These methods have been applied to civil infrastructures [XWYZ25], space frame structures [LXHL21], rebar detection [LKL21, LKS25], and landslide monitoring [ZCH23].
+In cases where prior information is outdated or unavailable, low-resolution preliminary scans are required, either indoors [QWTW21] or outdoors using UAVs [LLZ22].
+P4S optimizations aim to maximize coverage and overlap while minimizing the number of viewpoints [IGFER22, CSW25].
+Methods like CMA Evolution Strategy [RLGA22] and regular grid sampling [JL19, NB22] increase the likelihood of finding optimal viewpoints but also raise computational complexity.
+Jia’s approach using grid-based sampling [JL18, JL19, JL22] scores candidate viewpoints based on observed wall segments, though this approach is limited by its reliance on reference objects for registration.
+Noichl [NLB24] improve Jia’s method into 3D scene for an uniform covering by triangulating 3D scene into meshes and checking their visibility.
 For scenarios lacking BIM priors, structural elements can be surveyed and detected using low-cost sensors to serve as a preliminary prior for static station planning.
-A precise scan is then performed in a stop-and-go manner using a static scanner mounted on a robotic dog [9, 18].
-Our approach addresses these limitations by using a continuous Visibility Field (VF) model, which reduces optimization from 2D to 1D by focusing on structural convergence points like medial axis and joints, thus minimizing computation and enhancing adaptability.
+A precise scan is then performed in a stop-and-go manner using a static scanner mounted on a robotic dog [CKP25, HG25].
+Our approach addresses these limitations by using a continuous VF model, which reduces optimization from 2D to 1D by focusing on structural convergence points like medial axis and joints, thus minimizing computation and enhancing adaptability.
 2.2 Mobile-Based Exploration
-Coverage Path Planning (CPP) focuses on computing optimal, collision-free paths for robots to fully cover target areas [44].
-Techniques like Rapidly-exploring Random Trees (RRT) and their variants (RRT* [23], RRT# [37]) are widely used for autonomous exploration in unknown spaces, including UAVs [31] and wheeled robots [38].
-RRT-based methods maximize information gain while minimizing travel distance [17, 47, 53].
-Methods utilizing tensor fields for guiding robot movements enable efficient path routing in partially reconstructed indoor scenes [48, 46].
-Recent approaches like the Signed Distance Field (SDF) and Hamilton-Jacobi skeleton further improve path planning in known environments [36].
-These concepts share similarities with our visibility field, which leverages converging lines for efficient viewpoint planning.
-For known environments, algorithms like the receding horizon ”next-best-view” approach [3] and dual-resolution mapping schemes [6] balance detailed local mapping with efficient global exploration.
-For applications in autonomous driving and exploration, recent advances such as ”LookOut” [12] propose diverse multi-future prediction models that enhance autonomous navigation by predicting possible trajectories, balancing safety, and efficiency.
-3 Theoretical Foundations and Global Optimality
-This section formulates the Viewpoint Planning Problem (VPP) for static LiDAR scanning in a 2D polygonal environment and establishes the theoretical basis for restricting candidate viewpoints to the medial axis (MA).
-We first present the problem definition, including coverage and connectivity requirements, and then show that any line segment visible from an interior point is also visible from a corresponding set of MA points.
-This demonstrates that restricting the viewpoint search space to the MA incurs no loss of coverage capability.
-Finally, we prove that a minimal solution confined to the MA is globally optimal.
-3.1 Problem Definition
-Let $P$ be a 2D polygonal region (such as a floor plan) with possible obstacles or holes, while remaining topologically connected.
-The polygon’s boundary (including internal obstacles) is discretized into $n$ small line segments:
-$L=\{\,l_{1},\;l_{2},\;\dots,\;l_{n}\},$
-which represent the target structures (e.g., walls or windows) to be covered by static LiDAR scanners (hereafter referred to as viewpoints).
-Candidate Viewpoints
-We aim to position a set of viewpoints such that all segments in $L$ are fully visible, while the chosen viewpoints remain connected under an overlap criterion.
-Let $V=\{\,v_{1},\;v_{2},\;\dots,\;v_{m}\}$ be the set of candidate viewpoint locations.
-We will later establish that it is sufficient to confine $V$ to the medial axis of $P$, reducing a 2D search to a simpler 1D structure.
-Visibility
-we define a Completeness Table $C$, where $C_{ij}$ indicates whether viewpoint $V_{i}$ observes line segment $L_{j}$
-$C_{ij}=\begin{cases}1&\text{if }v_{i}\text{ has an unobstructed view of }l_{j}\\ 0&\text{otherwise}\end{cases}$ (1)
-Connectivity via Overlap
-Consider any pair of viewpoints $(v_{i},v_{k})$.
-Let $O_{ik}$ be their overlap ratio, i.e., the fraction of segments in $L$ that both $v_{i}$ and $v_{k}$ can simultaneously observe.
-For a threshold $\tau$, an edge $(v_{i},v_{k})$ exists in an overlap graph $G=(V,E)$ if $O_{ik}\geq\tau$.
-A chosen subset $S\subseteq V$ must then form a connected subgraph under this threshold.
-Mathematical Formulation
-We seek a minimal subset $S\subseteq V$ of viewpoints such that:
-  1. (a)
-    Coverage: Each segment $l_{j}\in L$ is observed by at least one viewpoint in $S$, i.e.,
-    $\forall\,l_{j}\in L,\;\exists\,v_{i}\in S:\;C_{ij}=1.$ (2)
-  2. (b)
-    Connectivity: The induced subgraph $(S,E_{S})$ is connected under the overlap threshold $\tau$, where
+Coverage Path Planning (CPP) focuses on computing optimal, collision-free paths for robots to fully cover target areas [TMMA21].
+Techniques like Rapidly-exploring Random Trees (RRT) and their variants (RRT* [KF11], RRT# [OF15]) are widely used for autonomous exploration in unknown spaces, including UAVs [MHS23] and wheeled robots [PSC23].
+RRT-based methods maximize information gain while minimizing travel distance [HNH18, XYH18, ZYX21].
+Methods utilizing tensor fields for guiding robot movements enable efficient path routing in partially reconstructed indoor scenes [XZY17, XZD24].
+Recent approaches like the Signed Distance Field (SDF) and Hamilton-Jacobi skeleton further improve path planning in known environments [NLMC23].
+These concepts share similarities with our VF, which leverages converging lines for efficient viewpoint planning.
+For known environments, algorithms like the receding horizon "next-best-view" approach [BKA16] and dual-resolution mapping schemes [CZR23] balance detailed local mapping with efficient global exploration.
+For applications in autonomous driving and exploration, recent advances such as "LookOut" [CCS21] propose diverse multi-future prediction models that enhance autonomous navigation by predicting possible trajectories, balancing safety, and efficiency.
+3 Problem Definition and Global Optimality
+We formalize the VPP for static LiDAR scanning across a collection of two-dimensional polygonal regions $\mathcal{P}=\{P_{1},P_{2},\dots,P_{k}\}$, each comprising structural boundaries and possible internal obstacles.
+The boundaries of these regions are uniformly discretized into a set of equal-length line segments $L=\{l_{1},l_{2},\dots,l_{n}\}$, representing the structural elements to be scanned.
+The objective is to select a set of viewpoints $V=\{v_{1},v_{2},\dots,v_{m}\}$ that satisfies two core constraints: coverage and connectivity.
+Coverage: Every segment $l_{j}\in L$ must be visible from at least one viewpoint $v_{i}\in V$.
+We formalize this visibility relationship using a binary coverage table $C$, where each entry $C_{ij}$ indicates whether $v_{i}$ observes $l_{j}$:
+$C_{ij}=\begin{cases}1&\text{if }v_{i}\text{ sees }l_{j},\\ 0&\text{otherwise}.\end{cases}$ (1)
+Connectivity: Let the overlap ratio $O_{ik}$ as the fraction of $L$ observed by both viewpoints $v_{i}$ and $v_{k}$.
+The induced subgraph $(S,E_{s})$ is connected under the overlap threshold $\tau$, where
 $E_{S}\;=\;\bigl\{(v_{i},v_{k})\in E\;\big|\;v_{i},v_{k}\in S,\;O_{ik}\geq\tau\bigl\}.$ (2)
-The optimization problem becomes:
+We aim to find the smallest subset $S\subseteq V$ such that:
 […]
-Since problem (4) generalizes the set cover problem (with additional connectivity constraints), it remains NP-hard.
-Nonetheless, by restricting $V$ to the medial axis (MA), we retain optimality while reducing the search complexity considerably.
-3.2 Skeleton Completeness and Visibility Propagation
-Medial Axis (MA)
-The MA of $P$ is the set of points inside $P$ that are equidistant to at least two boundary edges.
-In Blum’s grassfire transform [4], if the boundary of $P$ is ignited simultaneously, the inward-propagating flames collide and form the locus of the MA.
-Theorem 1 (Skeleton Completeness).
-Suppose an interior point $p\in P$ sees a set of boundary segments $L_{p}\subseteq L$. Then there exists a set of points $\{q_{l}\}\subseteq\mathrm{MA}(P)$ such that
-$L_{p}\;\;\subseteq\;\;\bigcup_{l\in L_{p}}\mathrm{Vis}(q_{l}).$
-Hence, for any interior coverage provided by $p$, there exists an MA-based set of points whose collective coverage reproduces that of $p$.
-Sketch of Proof.
-Using Blum’s “fire transformation,” the visibility from every segment propagates inward until it intersects the medial axis $\mathrm{MA}(P)$.
-Each line segment $l_{j}\in L_{p}$ must intersect $\mathrm{MA}(P)$ at some point $q_{j}$ visible to $l_{j}$.
-Thus,
-$\mathrm{Vis}(p)\;\;\subseteq\;\;\bigcup_{l_{j}\in\mathrm{Vis}(p)}\,\mathrm{Vis}(q_{j}).$
-Consequently, although a single point on the MA may not match $p$’s entire coverage, a suitable collection of MA points can, illustrated in 2.
-Restricting $V$ to $\mathrm{MA}(P)$ therefore does not diminish the achievable coverage.
-Figure 2: Illustration of skeleton completeness.
-(a) A simple polygon (black boundary) with its medial axis (gray arcs).
-(b) Coverage by an interior viewpoint $p$ can be replicated by a single skeletal viewpoint $q$ lying on the medial axis (MA).
-(c) In the general case, coverage from an interior viewpoint $p$ may require multiple skeletal viewpoints (here, $m$ and $n$) to fully reproduce $p$’s visibility.
-Thus, no coverage capability is lost by restricting candidate viewpoints to the MA.
-For clarity, boundary segments commonly observed by both $p$ and the MA viewpoints ($q$, or $m$ and $n$) are not drawn.
-3.3 Global Minimal Cover–Connectivity on the Skeleton
-By Theorem 1, any coverage by interior points can be equivalently transferred to a suitable set of points on the medial axis.
-Hence, in the formulation (4), restricting $V$ to $\mathrm{MA}(P)$ (including any branch or joint nodes) captures all feasible coverage solutions.
-Theorem 2 (Global Minimal Cover–Connectivity).
-Let $S^{*}\subseteq\mathrm{MA}(P)$ be a minimal solution to (4) that covers all $l_{j}\in L$ and satisfies the overlap-based connectivity requirement.
-Then $S^{*}$ is irreducible (i.e., removing any $v_{i}\in S^{*}$ breaks coverage or connectivity), and no smaller interior configuration can outperform $S^{*}$.
-Therefore, $S^{*}$ is globally optimal.
-Implication and Heuristic Search
-Exploiting the symmetry of the medial axis, no point outside $\mathrm{MA}(P)$ can provide superior coverage or connectivity.
-Consequently, a minimal cover–connectivity solution restricted to the skeleton cannot be improved by adding non-skeleton points.
-Hence, solving (4) directly on $\mathrm{MA}(P)$ yields a globally optimal arrangement in $P$.
-Although the problem remains NP-hard, standard polynomial-time greedy algorithm (e.g., [10]) often produce near-optimal results in practice.
-Once obtained, the minimal MA-based solution is irreducible, further affirming its global optimality for the entire polygon.
-Figure 3: Overview of the Viewpoint Network (VPN) optimization based on the Visibility Field (VF).
-(a) Input floor plan; (b) Visibility Field (VF); (c) Distance field of VF; (d) Converging lines and points; (e) Optimized Viewpoint Network.
+This NP-hard problem extends the classical set cover problem by incorporating additional connectivity constraints.
+We demonstrate that restricting $V$ to the medial axis (MA) preserves both complete coverage and network connectivity.
+As observation data in the VF naturally converge toward the MA and its joint points, placing viewpoints along these skeletal structures ensures comprehensive coverage.
+This property reduces VPN optimization from a 2D search to a 1D problem constrained to the MA.
+Consequently, a minimal cover on the MA remains globally optimal.
+Formal proofs of visibility completeness and global optimality for this reduction are provided in Appendix A.
 4 Method
 4.1 Overview
-Our proposed VF-Plan addresses the NP-hard complexity of the VPP through a greedy approach centered on the VF.
-Given a structural floor plan, VF-Plan constructs a Viewpoint Network (VPN) that achieves comprehensive coverage with minimal viewpoints, ensuring connectivity and registrability.
-Each viewpoint is selected to form a fully connected observation network, minimizing registration errors via Least Squares Fitting (LSF).
-As illustrated in Figure 3, the VF-Plan pipeline starts from a floor plan with structural elements (walls, columns, doors, windows) and proceeds through the steps of VF computation, distance field generation, identification of medial axis and joint points, and initialization of candidate viewpoints.
-A greedy algorithm then refines these candidates into a minimal, connected VPN.
-To optimize the computationally intensive task of determining visible wall segments from each viewpoint, we employ a BSP-tree[14] to accelerate processing.
-4.2 Visibility Field
+Our proposed method, VF-Plan, addresses the NP-hard complexity of the View Planning Problem (VPP) through a greedy strategy centered on the VF.
+Given a structural 2D floorplan, VF-Plan constructs a VPN that ensures full coverage with a minimal number of viewpoints, while maintaining connectivity and registrability.
+The selected viewpoints form a fully connected observation network, which can subsequently be used to minimize registration errors via Least Squares Fitting (LSF).
+As shown in fig.˜1, VF-Plan starts from a floorplan containing structural elements such as walls, columns, doors, and windows.
+The process includes VF computation, medial axis and joint-point extraction, and candidate viewpoint initialization.
+Overlap ratios between candidates are then evaluated, and a greedy algorithm selects the minimal subset forming a connected VPN.
+To accelerate the visibility checks for each viewpoint, we employ a BSP-tree [FKN80] with balanced splitting and no depth limit, starting from the median line at each iteration.
+Prior to BSP-tree construction and VF computation, the floorplan is preprocessed into closed polygons while preserving all structural elements.
+4.2 Visibility Field (VF)
 Laser scanning is limited to visible surfaces, leaving occluded areas unobserved.
-Therefore, visibility and coverage analysis are essential for viewpoint planning, with key objectives of maximizing observation completeness and registrability.
+Therefore, visibility and coverage analysis are essential for viewpoint planning, with key objectives of maximizing observation Coverage and registrability.
 The field of view is central to VPN optimization, defining the information captured at each viewpoint and guiding the selection of optimal viewpoints.
-A common method for evaluating the information scanned by a LiDAR sensor is based on the length of observed walls [22].
+A common method for evaluating the information scanned by a LiDAR sensor is based on the length of observed walls [JL22].
 Typically, wall length is counted either by ensuring both endpoints are visible or by partitioning walls into uniformly sampled segments within the scanner’s range.
@@
 Our approach instead uses the valid scanned angle to integrate relevant factors into a single index, uniformly scaled from 0 to $2\pi$.
-We define a static LiDAR scanner with a 360° field of view in the X-Y plane, a range from $r_{\text{min}}$ to $r_{\text{max}}$, and uniform scanning speed, as shown in Figure 4.
+We define a static LiDAR scanner with a 360° field of view in the X-Y plane, a range from $r_{\text{min}}$ to $r_{\text{max}}$, and uniform scanning speed, as shown in fig.˜2.
 Only objects within this range are observable.
@@
 Each viewpoint is assigned a View Angle, with the scene represented as a grid where each cell center is a viewpoint.
-Thus, the scene is represented as a Visibility Field, as shown in Figure 3(b), which aligns well with LiDAR properties and expert experience, displaying expected visibility patterns.
+Thus, the scene is represented as a VF, as shown in fig.˜3(b), which aligns well with LiDAR properties and expert experience, displaying expected visibility patterns.
+Figure 3: Visibility Field (VF) for static LiDAR scanning.
+(a, c) VF computed with only one wall (green highlight); (b, d) VF with all walls, revealing the global visibility structure.
+(a, b) show a rectangular room, and (c, d) depict a flat layout.
 4.3 Converging Lines and Joints
-Observation data naturally converge to the medial axis and then joints of the Visibility Field, so placing viewpoints along these skeletons and joints ensures full coverage.
-Thus, VPN optimization reduces from a 2D to a 1D problem along these skeletons.
-We use an medial axis algorithm [43] to identify skeletons and joint points, placing candidate viewpoints along these lines.
-Where overlap ratios are insufficient, additional viewpoints are added to ensure network cohesion.
-For any skeleton ridge $R_{ij}$ connecting joints $J_{i}$ and $J_{j}$, we calculate an Overlap Ratio $O_{ij}$ of the endpoints.
-When $O_{ij}$ falls below a threshold, a midpoint $J_{m}$ is added to split the ridge, ensuring connectivity.
-Both the skeleton joints and connecting points are called as converging points.
-And the resulting skeleton ridges are called converging lines.
-This iterative placement process continues until all converging points have adequate overlap for registrability.
+In the VF, observation data naturally converge toward the medial axis and its joint points.
+Placing viewpoints along these skeletal structures guarantees full coverage (section˜3, appendix˜A), reducing VPN optimization from a 2D search to a 1D problem.
+We extract the medial axis using [SP08] via OpenCV [Bra00] and detect joints by restricting inside room polygons: a skeleton point is marked as a joint candidate if at least three of its eight neighbors are also skeleton points.
+Connected-component analysis groups candidates, and each cluster’s centroid is retained as a representative joint.
+For each skeleton ridge $R_{ij}$ between joints $J_{i}$ and $J_{j}$, we compute the overlap ratio $O_{ij}$.
+If $O_{ij}$ is below the threshold, a midpoint $J_{m}$ is inserted to split the ridge, ensuring connectivity.
+We refer to all joints and added points as converging points, and their connecting ridges as converging lines.
+This process iterates until every converging point satisfies the overlap requirement for reliable registration.
 4.4 Overlap Ratio
-Figure 5: Overlap calculation between observed segments from different viewpoints, quantifying connectivity.
-The Overlap Ratio quantifies shared visibility between viewpoints, a requirement for robust registration.
-Five overlap ratios based on length and angle are defined, ensuring effective connectivity independent of distance alone.
-Given wall segments observed by viewpoint $V_{a}$ as $L_{a}$ with an observation angle $\theta_{a}$:
+Figure 4: Overlap computation between wall segments from two viewpoints. (a, b) Segments visible from viewpoints A and B. (c) Segments jointly visible, used to measure connectivity.
+The overlap ratio quantifies shared visibility between two viewpoints and is key for robust registration.
+We define five variants based on either segment length or angular alignment, evaluated for their ability to maintain connectivity independent of viewpoint spacing.
+As shown in fig.˜4, $L_{a}$ and $L_{b}$ denote wall segments visible from viewpoints $V_{a}$ and $V_{b}$, with observation angles $\theta_{a}$ and $\theta_{b}$:
 $\theta_{a}=\text{Angle}(V_{a},L_{a})$ (6)
@@
 $L_{ab}=L_{a}\cap L_{b}$ (7)
-The length-based overlap ratios are:
+Length-based ratios:
 $O^{\text{Min\_Len}}=\frac{L_{ab}}{\min(L_{a},L_{b})}$ (8)
@@
 $O^{\text{Union\_Len}}=\frac{L_{ab}}{L_{a}\cup L_{b}}$ (10)
-And the angle-based overlap ratios are:
+Angle-based ratios, where $\theta_{a}^{ab}$ and $\theta_{b}^{ab}$ are the angular spans of $L_{ab}$:
 $O^{\text{Union\_Ang}}=\frac{\theta_{a}^{ab}+\theta_{b}^{ab}}{\theta_{a}+\theta_{b}}$ (11)
@@
 4.5 Greedy Optimization for VPN Construction
-The VF-Plan greedy algorithm constructs an efficient Viewpoint Network (VPN) by starting with candidate viewpoints positioned strategically at key joints within the Visibility Field.
-For each pair of candidate viewpoints, an Overlap Ratio $O$ is calculated, with pairs showing adequate overlap retained as adjacencies in the VPN to ensure robust connectivity.
+The VF-Plan greedy algorithm constructs an efficient VPN by initializing candidate viewpoints at strategically selected joint points within the VF.
+For each pair of candidate viewpoints, an Overlap Ratio $O$ is computed to quantify shared visibility.
+Pairs exceeding a predefined overlap threshold are retained as edges in the VPN, ensuring robust connectivity.
+To enforce the coverage constraint, we construct a binary Coverage Table $C$ (see eq.˜1) that records whether each viewpoint observes each structural segment.
 Figure 5: Greedy Optimization of VPN.
 (a) Floorplan showing candidate viewpoints (labeled a to k) positioned along walls (numbered 0 to 17); (b) Connectivity graph, where edges link viewpoints with overlap ratio $O$ exceeding the threshold criterion; (c) Initial seed viewpoint selection based on maximum wall visibility (numbers near viewpoints indicate the count of observed walls); (d) - (f) Iterative expansion of the seed viewpoint set by adding adjacent viewpoints that provide the highest additional coverage, enhancing overall viewpoint arrangement.
-To systematically track visibility, we construct the Completeness Table $C$ (see Equation 1), which determines whether a given viewpoint $v_{i}$ observes a specific line segment $l_{j}$.
-To enhance precision in observation tracking, wall boundaries are uniformly partitioned into equal-length segments.
 Algorithm 1 Greedy Optimization of VPN
@@
   1. 1.
-    Initialize Completeness Table $C$
+    Initialize Coverage Table $C$
   2. 2.
@@
 return VPS as the optimized minimal VPN
-As outlined in Algorithm 1 and illustrated in Figure 6, the greedy optimization initiates by selecting the viewpoint with the highest coverage as the seed.
-The algorithm then iteratively expands by adding neighboring viewpoints that maximize additional coverage, ensuring each new viewpoint significantly contributes to covering line segments while constructing a minimal, fully connected Viewpoint Network (VPN).
+As outlined in Algorithm 1 and illustrated in fig.˜5, the greedy optimization initiates by selecting the viewpoint with the highest coverage as the seed.
+The algorithm then iteratively expands by adding neighboring viewpoints that maximize additional coverage, ensuring each new viewpoint significantly contributes to covering line segments while constructing a minimal, fully connected VPN.
 Upon completion of the initial search, the algorithm produces a single connected component encompassing all essential viewpoints, with sufficient overlap to facilitate stable connections.
@@
 Additionally, any node with more than three connecting edges from a cycle selects a Converging Joint to establish loops, thereby strengthening network connectivity.
-5 Experiments
 Figure 6: Performance of five proposed overlap ratios.
 The first row shows overlap ratios between moving and reference viewpoints; the second row illustrates room configurations, with black arrows indicating movement direction and green dots marking reference viewpoints.
-5.1 Dataset
+5 Experiments
 In this section, we evaluate the proposed VPN optimization algorithm across multiple scenarios and compare its performance with state-of-the-art methods in both indoor and outdoor environments.
@@
 This dataset provides a robust foundation for evaluating VPN optimization across a range of realistic settings.
+The pipeline is implemented in C++14 and compiled on Windows 11, running on an Intel Core i9 CPU (24 cores, 2.20 GHz) with 32 GB RAM.
+5.1 Dataset
 For indoor environments, we selected scenes with varied room layouts to evaluate the algorithm’s adaptability across multiple configurations.
-The Structure3D dataset [56] was included for its extensive collection of diverse Asian residential scenes, featuring both complex and compact room types frequently used in 3D modeling and synthetic data generation.
+The Structure3D dataset [ZZL20] was included for its extensive collection of diverse Asian residential scenes, featuring both complex and compact room types frequently used in 3D modeling and synthetic data generation.
 VPNs generated from this dataset can also facilitate synthetic LiDAR data production for advancing research in 3D scene understanding.
 Specific scenes such as 00007, 00009, 00071, 02970, and 02972, discussed in subsequent sections, are drawn from Structure3D.
-For further comparisons, we incorporated additional scenes from Zeng1 [51], Noichl [34], Xu [48], and TUB1 from the ISPRS benchmark [52, 24].
-For outdoor scenarios, we limited our approach to 2D site plans without considering terrain elevation, distinguishing our method from elevation-based approaches like [8].
-Selected scenes include the CCIT and Crowsnest datasets from the University of Calgary [20] and the ECUT dataset from East China University of Technology [8], which represent open-space environments suitable for VPN optimization.
+For further comparisons, we incorporated additional scenes from Zeng1 [ZLC22], Noichl [NB23], Xu [XZY17], and TUB1 from the ISPRS benchmark [ZZG24, KTA21].
+For outdoor scenarios, we limited our approach to 2D site plans without considering terrain elevation, distinguishing our method from elevation-based approaches like [CZH22].
+Selected scenes include the CCIT and Crowsnest datasets from the University of Calgary [JL18] and the ECUT dataset from East China University of Technology [CZH22], which represent open-space environments suitable for VPN optimization.
 To evaluate VPN quality, we rely on expert-labeled annotations as standard ground truth data is unavailable.
@@
 5.2 Performance Metrics
-The optimization algorithm aims to construct a minimal, fully connected viewpoint network with reduced redundancy.
+The optimization algorithm aims to construct a minimal, fully connected VPN with reduced redundancy.
 As most state-of-the-art methods achieve 100% coverage, we focus on two key metrics: the number of optimal viewpoints and network connectivity.
-Network connectivity is measured using the Weighted Average Path Length (WAPL), which reflects the compactness of the network [5].
+Network connectivity is measured using the Weighted Average Path Length (WAPL), which reflects the compactness of the network [BLM06].
 A shorter WAPL indicates stronger connectivity and a more efficient structure.
@@
 To ensure WAPL remains computable, we assign a large weight (100) to node pairs that lack a connecting path.
-5.3 Algorithm Parameter Effects
-5.3.1 Overlap Ratio
-We propose five overlap ratios according to the length of scanned walls and according angles to viewpoints in section 4.4.
-Here we evaluate them in two test indoor scenes, illustrated in Figure 7.
-Among them, $O^{\text{Min\_Len}}$ quickly reaches 100%, showing limited variability along movement paths, as seen from B to H in Figures 7(b) and (e), making it less suitable for dynamic scenes.
-The angle-based measures, $O^{\text{Union\_Ang}}$ and $O^{\text{Mean\_Ang}}$, exhibit non-monotonic behavior, particularly along paths from C to E, shown in Figures 7(c) and (f), which limits their usefulness for consistent registration.
+5.3 Ablation Study
+We evaluate overlap ratio type, scan radius, window inclusion, and, in appendix˜B, wall partition, VF grid resolution and overlap thresholds.
+VPN performance is largely insensitive to wall partition length and VF grid resolution, though these parameters affect computation time, while others require tuning for specific scenarios.
+Default settings for subsequent experiments are: $O^{\text{Mean\_Len}}$ for overlap computation, overlap ratio $0.4$ indoors and $0.3$ outdoors, $R_{\min}=0.6$ m indoors and $1.2$ m outdoors, $R_{\max}=30.0$ m indoors and $75.0$ m outdoors, wall partition $0.1$ m indoors and $1.0$ m outdoors, and VF grid resolution $0.02$ m indoors and $0.25$ m outdoors.
+5.3.1 Overlap Ratio Type
+We propose five overlap ratios according to the length of scanned walls and according angles to viewpoints in section˜4.4.
+Here we evaluate them in two test indoor scenes, illustrated in fig.˜6.
+Among them, $O^{\text{Min\_Len}}$ quickly reaches 100%, showing limited variability along movement paths, as seen from B to H in fig.˜6(b) and (e), making it less suitable for dynamic scenes.
+The angle-based measures, $O^{\text{Union\_Ang}}$ and $O^{\text{Mean\_Ang}}$, exhibit non-monotonic behavior, particularly along paths from C to E, shown in fig.˜6(c) and (f), which limits their usefulness for consistent registration.
 The $O^{\text{Union\_Len}}$ measure tends to be too small, often underestimating connectivity.
 Therefore, we select $O^{\text{Mean\_Len}}$ for subsequent experiments and analyses due to its balanced performance across different scenarios.
+Figure 7: Effect of $R_{\text{max}}$ on Viewpoint Placement and Network Efficiency.
+As $R_{\text{max}}$ increases from 15m to 75m, the valid placement area (where View Angle >> 1.0) grows, reducing the VC needed for optimized VPN coverage from 23 to 8.
 5.3.2 Scan Radius
@@
 For outdoor scenarios, we examine the impact of varying $R_{\text{max}}$, testing values at 15m, 30m, 45m, and 75m, with $R_{\text{min}}$ fixed at 1m.
-As illustrated in Figure 8, increasing $R_{\text{max}}$ expands the area of valid placements for viewpoints—defined where the View Angle exceeds 1.0—and reduces the Viewpoint Count (VC) needed for optimized coverage, from 23 to 14, 10, and 8.
+As illustrated in fig.˜7, increasing $R_{\text{max}}$ expands the area of valid placements for viewpoints—defined where the View Angle exceeds 1.0—and reduces the Viewpoint Count (VC) needed for optimized coverage, from 23 to 14, 10, and 8.
 Larger $R_{\text{max}}$ values are advantageous for open spaces, while smaller values require additional viewpoints for full coverage.
 These results indicate that our method aligns effectively with real-world scanner specifications and expert guidelines.
-Figure 8: Effect of $R_{\text{max}}$ on Viewpoint Placement and Network Efficiency.
-As $R_{\text{max}}$ increases from 15m to 75m, the valid placement area (where View Angle $>$ 1.0) grows, reducing the Viewpoint Count (VC) needed for optimized VPN coverage from 23 to 8.
-5.3.3 Wall Partition
-To improve station coverage accuracy, walls are partitioned into segments of varying lengths, as shown in Figure 9.
-We find that partitioning does not alter the final optimization result, but only affects the selection order of stations.
-Larger partition lengths (or no partitioning) place initial viewpoints at major intersections, optimizing computational cost and maintaining effective coverage.
-Figure 9: Effects of wall partitions. From left to right: unpartitioned walls and walls partitioned into segments of 1.0m, 0.5m, 0.1m, and 0.01m.
+Figure 8: Comparison of VPN structures with and without windows. Including windows (bottom row) adds complexity to the skeleton and slightly increases the VC.
 5.3.3 Effect of Windows
-Including windows in the model adds complexity to the skeleton structure, generating more detailed skeletons with additional joint points, as shown in Figure 10.
-When windows are considered, the VPN algorithm requires only a minimal increase in viewpoints to achieve complete coverage, as seen in Table 1.
-Without windows, the Viewpoint Count (VC) averages around 1.4 times the number of rooms, while with windows, it increases slightly to 1.6 times the room count.
-These results indicate that our algorithm effectively accommodates the added detail from windows.
-For precise window scans, including window frames in the optimization process is recommended.
-Figure 10: Comparison of VPN structures with and without windows. Including windows adds complexity to the skeleton and slightly increases the Viewpoint Count.
+Incorporating windows into the model increases the complexity of the skeleton structure by introducing additional joint points, resulting in a more detailed medial axis, as illustrated in fig.˜8.
+However, accurate scanning of windows is often essential in practical applications.
+To assess the impact of including windows, we evaluate the performance of our algorithm with and without windows considered in the optimization.
+When windows are included, the VPN algorithm requires only a slight increase in the number of viewpoints to achieve complete coverage, as shown in table˜1.
+On average, the number of viewpoints is approximately 1.4 times the number of rooms without windows, and increases modestly to 1.6 times with windows.
+These results demonstrate that the algorithm effectively handles the additional geometric detail introduced by windows.
+For applications requiring precise window coverage, including window frames in the optimization process is recommended.
 Scene 00007 00009 00071
@@
 VC (With Windows) 7 5 15
-Table 1: Viewpoint Count (VC) comparison for room coverage with and without windows across different scenes.
+Table 1: VC comparison for room coverage with and without windows across different scenes.
 5.4 Results and Analysis
-Scene WAPL VC Baseline Method
+[…]
 Expert Ours Baseline Expert Ours Baseline
-Zeng1 0.47 0.53 43.37 12 12 10 Zeng[51]
-Noichl 0.32 0.38 23.91 8 10 7 Noichl2023[34]
-TUB1 1.03 1.00 40.97 28 29 19 Zhai2024[52]
-Xu 0.35 0.31 28.69 8 8 7 Xu2017[48]
-CCIT 0.52 0.47 38.12 8 8 7 Jia2018[20]
-Crowsnest 0.50 0.49 34.02 9 8 8 Jia2018[20]
-ECUT 0.67 0.76 11.94 58 52 139 Chen2023[8]
-Table 2: Comparison of viewpoint counts (VC) and Weighted Average Path Length (WAPL) across indoor and outdoor scenes. Lower WAPL indicates more compact network connectivity.
+Zeng1 0.47 0.53 43.37 12 12 10 Zeng[ZLC22]
+Noichl 0.32 0.38 23.91 8 10 7 Noichl2023[NB23]
+TUB1 1.03 1.00 40.97 28 29 19 Zhai2024[ZZG24]
+Xu 0.35 0.31 28.69 8 8 7 Xu2017[XZY17]
+CCIT 0.52 0.47 38.12 8 8 7 Jia2018[JL18]
+Crowsnest 0.50 0.49 34.02 9 8 8 Jia2018[JL18]
+ECUT 0.67 0.76 11.94 58 52 139 Chen2023[CZH22]
+Table 2: Comparison of VC and WAPL across indoor and outdoor scenes. Lower WAPL indicates more compact network connectivity.
 We benchmarked our method against state-of-the-art approaches using datasets from prior studies, representing a range of indoor and outdoor scenes.
-Indoor datasets include Zeng1 [51], Noichl [34], and TUB1 from the ISPRS dataset [52].
-Outdoor datasets consist of CCIT and Crowsnest from the University of Calgary [20] and ECUT from East China University of Technology [8].
+Indoor datasets include Zeng1 [ZLC22], Noichl [NB23], and TUB1 from the ISPRS dataset [ZZG24].
+Outdoor datasets consist of CCIT and Crowsnest from the University of Calgary [JL18] and ECUT from East China University of Technology [CZH22].
 Indoor Results.
-As shown in Figure 12, our method’s viewpoint optimization closely aligns with expert-designed networks, achieving comprehensive coverage and creating compact, interconnected structures.
+As shown in fig.˜9, our method’s viewpoint optimization closely aligns with expert-designed networks, achieving comprehensive coverage and creating compact, interconnected structures.
 This optimized network design enhances registration accuracy and connectivity, even in complex indoor settings.
 Outdoor Results.
-For outdoor scenes, our method effectively adapts to open spaces, as illustrated in Figure 13.
+For outdoor scenes, our method effectively adapts to open spaces, as illustrated in fig.˜10.
 Adjustments to the $R_{\text{max}}$ parameter notably improved performance in expansive environments, reducing the required number of viewpoints without compromising coverage.
-Table 2 demonstrates that our method performs comparably or better than expert-designed networks, achieving similar viewpoint counts (VC) and lower Weighted Average Path Length (WAPL).
-This outcome indicates a more compact and cohesive network structure.
-Unlike most baselines that often yield fragmented networks (highlighted by green ellipses in Figures 12 and 13), our approach maintains low WAPL scores, which reflect enhanced connectivity and minimized redundancy.
-In indoor environments, our method achieves a 98% reduction in WAPL compared to baseline methods, creating highly connected networks well-suited for confined spaces.
-For complex outdoor scenarios, such as ECUT, our approach achieves a 91% WAPL reduction over baselines, demonstrating scalability and adaptability across both intricate indoor layouts and expansive outdoor scenes.
-To further validate the optimized viewpoints, we simulated static LiDAR scanning at each optimized viewpoint using Helios++ [13].
-As shown in Figure 11, the simulated LiDAR point clouds confirm complete scene coverage, with effective overlap between adjacent scans across both indoor and outdoor environments.
-Figure 11: Simulated LiDAR points from optimal viewpoints. Each scan is uniquely colored. Top row shows top views, and bottom row shows oblique views with tripods indicating viewpoints.
-Figure 12: Viewpoint planning results for indoor environments.
-Red dots represent individual viewpoints, while green ellipses indicate clusters of viewpoints that are not connected to the main network.
+Table˜2 shows that our method matches or surpasses expert-designed networks while significantly outperforming baseline methods.
+Across all scenes, our VC remains comparable to experts, differing by at most one viewpoint, yet our WAPL is consistently lower in five of seven cases, indicating more compact and cohesive networks.
+Compared to baselines, the improvement is substantial: in indoor scenes such as Zeng1 and Noichl, WAPL is reduced by 98% and 97% respectively, eliminating the fragmented connectivity seen in baseline outputs (green ellipses in figs.˜9 and 10).
+In challenging outdoor cases, reductions remain high, including 91% in ECUT and 87% in CCIT, demonstrating scalability to large and complex environments.
+These results confirm that our VF-based optimization preserves expert-level coverage while achieving markedly superior network compactness over existing automated methods.
+To further validate the optimized viewpoints, we simulated static LiDAR scanning at each optimized viewpoint using Helios++ [EYW22].
+As shown in fig.˜11, the simulated LiDAR point clouds confirm complete scene coverage, with effective overlap between adjacent scans across both indoor and outdoor environments.
+Figure 9: Viewpoint planning results for indoor environments. Red dots represent individual viewpoints, while green ellipses indicate clusters of viewpoints that are not connected to the main network.
 Figure 10: Viewpoint planning results for outdoor environments.
 Red dots represent individual viewpoints, while green ellipses indicate clusters of viewpoints that are not connected to the main network.
+Figure 11: Simulated LiDAR points from optimal viewpoints. Each scan is uniquely colored. Top row shows top views, and bottom row shows oblique views with tripods indicating viewpoints.
 6 Conclusions
-We introduced a novel solution to the Viewpoint Planning Problem and Art Gallery Problem in static LiDAR scanning by leveraging a Visibility Field to reduce optimization complexity from 2D to 1D.
-This VF-based approach enables the creation of a minimal, fully connected viewpoint network that ensures comprehensive coverage with reduced redundancy, advancing viewpoint planning for static LiDAR applications.
+We presented a novel solution to the VPP and the AGP in static LiDAR scanning on 2D floor plans, leveraging the VF to reduce the optimization complexity from two dimensions to one.
+This VF-based approach enables the construction of a minimal, fully connected VPN that ensures complete coverage with reduced redundancy, advancing viewpoint planning for static LiDAR applications such as BIM construction.
 Our method offers a scalable and efficient framework applicable in structured and semi-structured environments, addressing real-world challenges in 3D data collection and autonomous navigation.
 Our VF framework adapts well to real-world settings with potential obstructions by integrating partial occlusion handling based on initial scans.
-In summary, this method provides a practical and robust solution for static viewpoint planning, with promising extensions to dynamic environments.
-Future work will focus on its application for large-scale LiDAR data collection and synthetic data generation, and expanding the method to support path planning in both 2D and 3D scenarios for more versatile deployment in complex environments.
-CRediT authorship contribution statement
-Biao Xiong: Writing - Original draft, Validation, Methodology, Investigation, Project administration.
-Longjun Zhang: Original draft, Validation, Software, Methodology.
-Ruiqi Huang: Validation, Software, Methodology, Data curation.
-Junwei Zhou: Validation, Software, review & editing.
-Bojian Wu: Writing - review & editing, Writing - original draft, Visualization.
-Fashuai Li: Writing - review & editing, Supervision, Resources, Conceptualization.
-Declaration of competing interest
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+The current formulation assumes accurate and complete 2D floor plans, which may not be available in all scenarios.
+In cases of noisy input or partial scans, the quality of the VF and the resulting VPN may degrade due to inaccuracies in boundary detection or incomplete visibility information.
+While preliminary layouts can be generated using low-cost SLAM systems or coarse surveys, further robustness to noise, clutter, and missing geometry remains an open challenge.
+Additionally, the current implementation focuses on planar (2D) environments; extending the approach to handle 3D structures with significant vertical complexity will require additional algorithmic considerations.
+Future work will explore its application to large-scale LiDAR data collection and synthetic data generation, as well as integration with path planning in both 2D and 3D scenarios for more versatile deployment in complex environments.
+Supplementary Materials
+Appendix A Theoretical Foundations, Medial Axis Convergence, and Greedy Optimality
+The key idea of VF-Plan is that visibility information concentrates on a one dimensional skeleton of the polygonal domain.
+In the visibility field, visibility propagates from the boundary and accumulates on the medial axis, which reduces the search from two dimensions to one.
+We formalize this convergence, prove that restricting candidate viewpoints to the medial axis preserves global optimality, and present quantitative performance guarantees for the greedy selection strategy.
+A.1 Skeleton Completeness and Visibility Propagation
+Medial Axis (MA).
+The medial axis of a polygon $P$ is the locus of points that are equidistant to at least two boundary elements.
+In Blum’s grassfire transform [Blu67] the boundary ignites at time zero and inward fronts meet on the medial axis, which explains the concentration of visibility information on the skeleton.
+Theorem 1 (Skeleton Completeness)
+Let $p\in P$ see a set of boundary segments $L_{p}\subseteq L$. There exists a finite set $\{q_{\ell}\}\subseteq\mathrm{MA}(P)$ such that
+$L_{p}\subseteq\bigcup_{\ell\in L_{p}}\mathrm{Vis}(q_{\ell}).$ (15)
+Hence any interior coverage provided by $p$ can be reproduced by viewpoints on the medial axis.
+Sketch of Proof.
+By the grassfire view [Blu67] visibility from each boundary segment propagates inward until it meets $\mathrm{MA}(P)$.
+For every $l_{j}\in L_{p}$ there exists $q_{j}\in\mathrm{MA}(P)$ that sees $l_{j}$.
+Therefore $\mathrm{Vis}(p)\subseteq\bigcup_{l_{j}\in\mathrm{Vis}(p)}\mathrm{Vis}(q_{j})$.
+Figure 12 illustrates the construction.
+Restricting candidates to $\mathrm{MA}(P)$ does not reduce achievable coverage.
+A.2 Global Minimal Cover and Connectivity on the Skeleton
+By theorem˜1 any feasible interior solution transfers to the medial axis. In the eq.˜3 restricting $V$ to $\mathrm{MA}(P)$ including branch and joint nodes therefore captures all feasible solutions.
+Theorem 2 (Global Minimal Cover and Connectivity)
+Let $S^{*}\subseteq\mathrm{MA}(P)$ be a minimal solution to eq.˜3 that covers all $l_{j}\in L$ and satisfies the required overlap–based connectivity.
+Then $S^{*}$ is irreducible and no smaller configuration in $P$ can achieve the same objectives.
+Therefore $S^{*}$ is globally optimal.
+Implication and heuristic search.
+No point outside $\mathrm{MA}(P)$ improves coverage or connectivity over a minimal skeletal solution.
+Solving eq.˜3 on $\mathrm{MA}(P)$ yields a configuration that is optimal for the whole polygon.
+The problem remains NP–hard, yet a greedy strategy is effective in practice and admits quantitative guarantees for the number of viewpoints.
+A.3 Greedy Selection and Suboptimality Bounds
+Let $n=|L|$ be the number of discretized boundary segments.
+Ignoring connectivity, selecting viewpoints to cover $L$ is a set cover instance.
+The classic greedy rule achieves an approximation factor $H_{n}\leq 1+\ln n$ for the number of selected sets [Chv79].
+Our solver first runs greedy coverage on medial–axis candidates and then augments connectivity on the induced viewpoint graph whose edge weights are $w_{ij}=1-O_{ij}$ where $O_{ij}$ is the overlap ratio.
+Connecting the greedy terminals with a Steiner tree in this graph adds only connector viewpoints along medial–axis paths.
+Using a constant–factor approximation for Steiner tree on graphs, the number of added connectors is within a constant factor of the optimum connector count.
+Combining the two stages gives
+$|S_{\text{greedy}}|\;\leq\;H_{n}\cdot\mathrm{OPT}_{\text{cover}}\;+\;\alpha\,\mathrm{OPT}_{\text{conn}},$ (16)
+where $\alpha$ is a constant and $\mathrm{OPT}_{\text{cover}}$ and $\mathrm{OPT}_{\text{conn}}$ are the minimal numbers of coverage and connector viewpoints on $\mathrm{MA}(P)$.
+When connectivity is not stricter than coverage, $\mathrm{OPT}_{\text{conn}}\leq\mathrm{OPT}_{\text{cover}}$, which yields the compact bound
+$|S_{\text{greedy}}|\;\leq\;(H_{n}+\alpha)\,\mathrm{OPT},$ (17)
+with $\mathrm{OPT}$ the minimal skeletal solution of eq.˜3.
+Empirically our solutions match or improve expert VC in five of seven scenes and reduce WAPL by up to $98\%$ relative to automated baselines while remaining within one viewpoint of expert designs in most cases, which corroborates the tightness of the bound in practice.
+Figure 12: Illustration of skeleton completeness.
+(a) A simple polygon (black boundary) with its medial axis (gray arcs).
+(b) Coverage by an interior viewpoint $p$ can be replicated by a single skeletal viewpoint $q$ lying on the medial axis (MA).
+(c) In the general case, coverage from an interior viewpoint $p$ may require multiple skeletal viewpoints (here, $m$ and $n$) to fully reproduce $p$’s visibility.
+Thus, no coverage capability is lost by restricting candidate viewpoints to the MA.
+For clarity, boundary segments commonly observed by both $p$ and the MA viewpoints ($q$, or $m$ and $n$) are not drawn.
+Appendix B More Ablation Study
+B.0.1 Effect of Wall Partition
+To evaluate the impact of wall partition length, wall boundaries are divided into segments of varying lengths including 1.0 m, 0.5 m, 0.1 m, and 0.01 m, as illustrated in fig.˜13.
+Experimental results show that the partition length does not affect the final optimization outcome, due to the continuous nature of the VF employed in our method.
+The resolution of partitioning influences only the order in which viewpoints are selected.
+Coarser partitions, corresponding to longer segments or no partitioning, tend to place initial viewpoints at major structural intersections.
+A coarse partition reduces computational cost while maintaining effective coverage.
+Based on this observation, we set the default partition length to 0.1 m for indoor scenes and 1.0 m for outdoor scenes in all subsequent experiments.
+Figure 13: Effects of wall partitions. From left to right: unpartitioned walls and walls partitioned into segments of 1.0m, 0.5m, 0.1m, and 0.01m.
+B.1 Effect of Overlap Ratio Threshold
+We analyze the influence of overlap ratio (OR) on VPN optimization using the indoor Scene_00001 dataset.
+The OR defines the minimum visible segment proportion shared between two viewpoints for them to be considered connected.
+Experiments were conducted for OR values ranging from $0.3$ to $0.7$, with all other parameters fixed: $R_{\min}=0.6$ m, $R_{\max}=30.0$ m, resolution $0.02$ m, and wall partition length $0.2$ m.
+Results in tables˜3, 14 and 15 show that increasing OR generally enhances network connectivity but often requires more viewpoints to maintain full coverage.
+For example, raising OR from $0.3$ to $0.7$ increases the number of viewpoints from $10$ to $18$ (an 80% increase) while WAPL rises from $0.38$ to $0.44$.
+Moderate OR values, such as $0.4$, achieve a balanced trade-off with only $9$ viewpoints, a WAPL of $0.40$, and stable runtime ($\approx 0.74$ s across all settings).
+Very high OR values ($\geq 0.6$) lead to sharp viewpoint growth without proportional gains in WAPL, indicating diminishing returns in connectivity improvement.
+These observations suggest that an OR in the range 0.30.30.50.5 is sufficient for most scanning tasks, providing robust connectivity while keeping network size compact.
+OR 0.3 0.4 0.5 0.6 0.7
+VC 10 9 11 13 18
+WAPL 0.38 0.40 0.45 0.45 0.44
+Time (s) 0.72 0.75 0.73 0.74 0.74
+Table 3: Effect of overlap ratio (OR) on Scene_00001.
+[…]
+Figure 14: Effect of overlap ratio (OR) on Scene_00001. WAPL and runtime are multiplied by 10 for scale comparability.
+Figure 15: Effect of overlap ratio (OR) on viewpoint planning in Scene_00001.
+Increasing OR from 0.3 to 0.7 improves connectivity but raises viewpoint count (VC), showing a trade-off between coverage and compactness, while WAPL and runtime remain stable.
+B.2 Effect of VF Grid Resolution
+We examine the impact of VF grid resolution on VPN optimization using the indoor Zeng_SceneA and outdoor CCIT datasets.
+For Zeng_SceneA, parameters are $R_{\min}=0.6$ m, $R_{\max}=30.0$ m, overlap ratio $0.4$, and wall partition $0.2$ m.
+For CCIT, parameters are $R_{\min}=1.2$ m, $R_{\max}=75.0$ m, overlap ratio $0.3$, and wall partition $1.0$ m.
+All other settings remain fixed, and only the grid resolution is varied.
+As shown in tables˜5, 16, 17 and 18, resolution changes have minimal influence on VPN count or WAPL, with variations under $5\%$ for both datasets.
+In Zeng_SceneA, VPN count ranges from $10$ to $12$ and WAPL from $0.52$ to $0.46$ as resolution decreases from $0.01$ m to $0.05$ m, while runtime drops from $1.57$ min to $0.03$ min.
+In CCIT, VPN count varies from $8$ to $9$ and WAPL from $0.43$ to $0.44$ as resolution decreases from $0.15$ m to $0.35$ m, with runtime falling from $0.21$ min to $0.03$ min.
+Runtime scales approximately quadratically with grid resolution due to the growth in visibility evaluations.
+These results confirm that the VF-based approach depends primarily on the topological structure of the floorplan, allowing the use of lower resolutions that capture necessary geometric detail while significantly improving runtime.
+Table 4: Effect of resolution on Viewpoint count, WAPL, and runtime for indoor (Zeng_SceneA) and outdoor (Outdoor_CCIT) cases.
+Zeng_SceneA Outdoor_CCIT
+Resolution (m) 0.01 0.02 0.03 0.04 0.05 0.15 0.20 0.25 0.30 0.35
+VC 11 10 11 11 12 8 8 8 8 9
+WAPL 0.52 0.52 0.49 0.48 0.46 0.43 0.45 0.43 0.43 0.44
+Time (min) 1.57 0.16 0.07 0.04 0.03 0.21 0.08 0.05 0.03 0.03
+[…]
+Figure 16: Effect of resolution on Viewpoint count (VC), WAPL, and runtime for indoor (Zeng_SceneA).
+[…]
+Figure 17: Effect of resolution on Viewpoint count (VC), WAPL, and runtime for outdoor (CCIT).
+Figure 18: Effect of resolution on VF-based viewpoint planning.
+The top row shows results for the indoor Zeng_SceneA case with resolutions ranging from 0.01 m to 0.05 m, and the bottom row shows results for the outdoor CCIT case with resolutions ranging from 0.15 m to 0.35 m.
+The VF’s topological nature results in low sensitivity to resolution changes in both scenarios.
+B.3 Time and Space Complexity Analysis
+We evaluate the computational complexity of VF-based VPN optimization in a large-scale outdoor town scenario.
+To assess scalability, the original dataset is cropped into subregions with approximately $200,400,600,800,$ and $1000$ viewpoints (VC).
+The full town comprises 637 buildings within a block of $976\,\text{m}\times 893\,\text{m}$, totaling $80{,}549\,\text{m}^{2}$ of built area with an average footprint of $126\,\text{m}^{2}$.
+All parameters are fixed across experiments: VF resolution $2.0\,\text{m}$, minimum scan radius $0.6\,\text{m}$, maximum scan radius $600\,\text{m}$, wall density $5\,\text{m}$, and overlap ratio $0.3$.
+This setup enables controlled scaling of VC while preserving realistic geometric complexity (tables˜5, 19 and 20).
+Table˜5 shows that memory usage grows nearly linearly with VC, from $265\,\text{MB}$ at VC=200 to $5703\,\text{MB}$ at VC=1000 (a 21.5× increase).
+On a logarithmic scale, $\log_{10}(\text{RAM})$ increases by only 1.34 across the five settings, confirming mild scaling.
+By contrast, runtime grows superlinearly: from $2.68\,\text{min}$ at VC=200 to $1505.36\,\text{min}$ ($\approx 25\,\text{hrs}$) at VC=1000, corresponding to a 561× increase.
+The logarithmic slope of $\log_{10}(\text{Runtime})$ averages 0.55 per VC-doubling, consistent with near-quadratic growth due to pairwise visibility evaluations.
+Comparative scaling further highlights runtime as the primary bottleneck.
+For example, increasing VC from 400 to 800 multiplies memory by 3.2× (1018 to 3268 MB) but runtime by 31× (30.8 to 967.5 min).
+Despite this, memory remains within typical workstation limits (<6 GB), while runtime quickly becomes prohibitive at VC $\geq 1000$.
+These findings demonstrate that VF-based optimization is tractable at city-block scales but requires algorithmic acceleration—such as hierarchical visibility pruning, parallel BSP-tree queries, GPU acceleration, or distributed processing—for deployment in very large regions.
+[…]
+Figure 19: Log-scaled memory and runtime vs.
+viewpoint count (VC) for the town dataset.
+Both $\log_{10}$(RAM/MB) and $\log_{10}$(Runtime/min) increase with CV, highlighting time and space complexity trends.
+Table 5: Time and space complexity analysis for the town dataset with cropped regions.
+VC
+VC
+RAM
+$[MB]$
+$\log_{10}$
+(RAM/MB)
+Runtime
+$[min]$
+$\log_{10}$
+(Runtime/min)
+Building
+Count
+Block Area
+$m\times m]$
+Total Building
+Area $m^{2}]$
+Avg Area
+Building $m^{2}]$
+200 197 265 2.42 2.68 0.43 132 $520\times 292$ 22059 167
+400 409 1018 3.01 30.84 1.49 274 $965\times 366$ 39359 144
+600 553 1683 3.23 71.02 1.85 379 $976\times 441$ 49428 130
+800 757 3268 3.51 967.48 2.99 529 $976\times 588$ 67448 128
+1000 954 5703 3.76 1505.36 3.18 637 $976\times 893$ 80549 126
+Figure 20: Optimized VPNs for the town dataset. Panels (a) to (e) show cropped regions for viewpoint counts $\mathrm{VC}=200,400,600,800,1000$ respectively.